# Connectedness of level sets for non-degenerate integrable systems that extend complexity one torus actions

### v2

@@ 6. Non-degenerate singular points
Proof.
For simplicity, assume that $g(p)=0$. By the Marle-Guillemin-Sternberg local normal form theorem, we may assume that $M$ is a local model $Y=T\times_{H}{\mathfrak{h}}^{\circ}\times{\mathbb{C}}^{h+1}$, that $\Phi=\Phi_{Y}$, and that $p=[1,0,0]$. Let $j\colon{\mathbb{C}}^{h+1}\to T\times_{H}{\mathfrak{h}}^{\circ}\times{\mathbb{C}}^{h+1}$ denote the inclusion $z\mapsto[1,0,z]$. By Lemma 3.2, $p$ is a critical point of $g$ modulo $\Phi$. Hence, the $(T\times{\mathbb{R}})$-orbit $\mathcal{O}$ through $p$ coincides with the $T$-orbit $\mathfrak{O}$ through $p$ and the Taylor polynomial $T^{1}_{0}(j^{*}g)$ of degree one at $0\in{\mathbb{C}}^{h+1}$ vanishes. Thus we may identify $(T_{p}\mathcal{O})^{\omega}/T_{p}\mathcal{O}=(T_{p}\mathfrak{O})^{\omega}/T_{p}\mathfrak{O}$ with ${\mathbb{C}}^{h+1}$ so that the homogeneous moment map for the linearized action at $p$ is the Taylor polynomial of degree two $T^{2}_{0}(j^{*}f)=(T^{2}_{0}(j^{*}\Phi),T^{2}_{0}(j^{*}g))=(\Phi_{H},T^{2}_{0}(j^{*}g))$, where $\Phi_{H}\colon{\mathbb{C}}^{h+1}\to{\mathfrak{h}}^{*}$ is the homogeneous moment map for the $H$-action. .
Proposition 6.4.
Let $\left(M,\omega,{\mathfrak{t}}\times{\mathbb{R}},f:=\left(\Phi,g\right)\right)$ be an integrable system such that $\left(M,\omega,\Phi\right)$ is a complexity one $T$-space. Assume that $p\in\Phi^{-1}(0)\cap g^{-1}(0)$ is tall and non-degenerate but does not have purely elliptic type. If the defining polynomial of the local model for $p$ has degree 1, then there exist an open neighborhood $U$ of $[p]$ in the reduced space ${M/\!/T}$ and a map $\Psi\colon U\to{\mathbb{R}}^{2}$ taking $[p]$ to $(0,0)$ satisfying:
  1. (a)
    $\Psi\colon U\to\Psi(U)$ is a diffeomorphism onto an open set; and
  2. (b)
    $\overline{g}\circ\Psi^{-1}(x,y)=x^{2}-y^{2}$ for all $(x,y)\in\Psi(U)$.
Moreover, $p$ has one hyperbolic block and connected $T$-stabilizer; every other block is elliptic.
Proof.
By the Marle-Guillemin-Sternberg local normal form theorem, we may assume that $M$ is a tall local model $Y=T\times_{H}{\mathfrak{h}}^{\circ}\times{\mathbb{C}}^{h+1}$, that $\Phi=\Phi_{Y}$, and that $p=[1,0,0]$. As in the proof of Lemma 6.3, by Lemma 3.2 we may identify $(T_{p}\mathcal{O})^{\omega}/T_{p}\mathcal{O}$ with ${\mathbb{C}}^{h+1}$ so that the homogeneous moment map for the linearized action at $p$ is the Taylor polynomial of degree two $(\Phi_{H},T^{2}_{0}(j^{*}g))$.
Since the defining polynomial $P\colon{\mathbb{C}}^{h+1}\to{\mathbb{C}}$ has degree 1, we may assume that $P(z)=z_{0}$. Hence, ${\mathbb{C}}^{h+1}\simeq{\mathbb{C}}\times{\mathbb{C}}^{h}$ and $H\simeq(S^{1})^{h}$, where $(S^{1})^{h}$ acts on ${\mathbb{C}}^{h}$ in the standard fashion and trivially on ${\mathbb{C}}$. Thus the span of $\{\Phi_{H}^{\xi}\mid\xi\in{\mathfrak{h}}\}$ equals the span of $\{|z_{k}|^{2}\mid k\geq 1\}$. Moreover, since the Taylor polynomial $T^{2}_{0}(j^{*}g)$ is $H$-invariant, by Remark 2.12, it lies in the span of $\{{\mathfrak{R}}(z_{0}^{2}),{\mathfrak{I}}(z_{0}^{2}),|z_{i}|^{2}\colon i=0,\ldots,h\}$. Therefore there exist constants $a_{x},a_{y},a_{\rho}\in{\mathbb{R}}$ and $\xi\in{\mathfrak{h}}$ such that $T^{2}_{0}(j^{*}g)$ is given by
(6.5) $a_{x}{\mathfrak{R}}(z_{0}^{2})+a_{y}{\mathfrak{I}}(z_{0}^{2})+a_{\rho}|z_{0}|^{2}+\Phi_{H}^{\xi}.$
$a_{x}{\mathfrak{R}}(z^{2}_{0})+a_{y}{\mathfrak{I}}(z^{2}_{0})+a_{\rho}|z_{0}|^{2}+\sum\limits_{k=1}^{h}\lambda_{k}|z_{k}|^{2},$
where $\lambda_{k}\in{\mathbb{R}}\smallsetminus\{0\}$ for $k=1,\ldots,h$. The eigenvalues of the corresponding symplectic endomorphism are given by
±$\pm 2\sqrt{a_{x}^{2}+a_{y}^{2}-a_{\rho}^{2}},\,\,\pm 2\lambda_{1}\sqrt{-1},\,\,\ldots,\,\,\pm 2\lambda_{h}\sqrt{-1}.$
Hence, since $p$ is non-degenerate and does not have purely elliptic type, Remark 6.1 implies that $a_{x}^{2}+a_{y}^{2}-a_{\rho}^{2}>0$, that $p$ has one hyperbolic block, and that the remaining blocks are all elliptic.
By Lemma 2.8, the inclusion $i\colon{\mathbb{C}}\to Y$ sending $z_{0}$ to $[1,0,(z_{0},0,\ldots,0)]$ induces a diffeomorphism $\overline{i}\colon{\mathbb{C}}\to{Y/\!/T}$. By definition of the inclusion $j$ and by (6.5),
$T^{2}_{0}(\overline{i}^{*}\overline{g})(z_{0})=a_{x}{\mathfrak{R}}(z_{0}^{2})+a_{y}{\mathfrak{I}}(z_{0}^{2})+a_{\rho}|z_{0}|^{2},$
and so 00 is a non-degenerate critical point of $\overline{i}^{*}\overline{g}$ of index 1. Hence, by the Morse lemma, there exists an open neighborhood $V$ of $(0,0)\in{\mathbb{R}}^{2}$ and a map $\psi\colon V\to{\mathbb{C}}$ taking $(0,0)$ to 00 that is a diffeomorphism onto an open set such that $(\overline{i}\circ\psi)^{*}\overline{g}(x,y)=x^{2}-y^{2}$ for all $(x,y)\in V$. Setting $U:=(\overline{i}\circ\psi)(V)$ and $\Psi\colon U\to{\mathbb{R}}^{2}$ to be $(\overline{i}\circ\psi)^{-1},$ the result follows. ∎
Proposition 6.6.
Let $\left(M,\omega,{\mathfrak{t}}\times{\mathbb{R}},f:=\left(\Phi,g\right)\right)$ be an integrable system such that $\left(M,\omega,\Phi\right)$ is a complexity one $T$-space. Assume that $p\in\Phi^{-1}(0)\cap g^{-1}(0)$ is tall and non-degenerate but does not have purely elliptic type. If the defining polynomial of the local model for $p$ has degree 2, then there exist an open neighborhood $U$ of $[p]$ in the reduced space ${M/\!/T}$ and a map $\Psi\colon U\to{\mathbb{R}}^{2}$ taking $[p]$ to $(0,0)$ satisfying:
  1. (1)
    $\Psi\colon U\to\Psi(U)$ is a homeomorphism onto an open set;
  2. (2)
    $\Psi$ restricts to a diffeomorphism on $U\smallsetminus\{[p]\}$; and
  3. (3)
    $(\overline{g}\circ\Psi^{-1})(x,y)=y$ for all $(x,y)\in\Psi(U)$.
Moreover, either $p$ has one focus-focus block, or it has one hyperbolic block and disconnected $T$-stabilizer; in either case, every other block is elliptic.
Proof.
By the Marle-Guillemin-Sternberg local normal form theorem, we may assume that $M$ is a tall local model $Y=T\times_{H}{\mathfrak{h}}^{\circ}\times{\mathbb{C}}^{h+1}$, that $\Phi=\Phi_{Y}$, and that $p=[1,0,0]$. As in the proof of Lemma 6.3, by Lemma 3.2 we may identify $(T_{p}\mathcal{O})^{\omega}/T_{p}\mathcal{O}$ with ${\mathbb{C}}^{h+1}$ so that the homogeneous moment map for the linearized action at $p$ is the Taylor polynomial of degree two $(\Phi_{H},T^{2}_{0}(j^{*}g))$. Since the defining polynomial $P\colon{\mathbb{C}}^{h+1}\to{\mathbb{C}}$ has degree 2, we may assume that we are in one of the following two cases (cf. Examples 2.5 and 2.6).
Case 1: $\mathbf{P(z)=z_{0}^{2}}$. In this case, ${\mathbb{C}}^{h+1}\simeq{\mathbb{C}}\times{\mathbb{C}}^{h}$ and $H\simeq\mathbb{Z}_{2}\times(S^{1})^{h}$, where $\mathbb{Z}_{2}$ acts on ${\mathbb{C}}$ by multiplication with $-1$ and trivially on ${\mathbb{C}}^{h}$, and $(S^{1})^{h}$ acts on ${\mathbb{C}}^{h}$ in the standard fashion and trivially on ${\mathbb{C}}$. Thus the span of $\{\Phi_{H}^{\xi}\mid\xi\in{\mathfrak{h}}\}$ equals the span of $\{|z_{k}|^{2}\mid k\geq 1\}$. Moreover, since the Taylor polynomial $T^{2}_{0}(j^{*}g)$ is $H$-invariant, by Remark 2.12, it lies in the span of $\{{\mathfrak{R}}(z_{0}^{2}),{\mathfrak{I}}(z_{0}^{2}),|z_{i}|^{2}\colon i=0,\ldots,h\}$. As in the proof of Proposition 6.4, this implies that there exist constants $a_{x},a_{y},a_{\rho}\in{\mathbb{R}}$ satisfying $a_{x}^{2}+a_{y}^{2}-a_{\rho}^{2}>0$ and $\xi\in{\mathfrak{h}}$ such that $T^{2}_{0}(j^{*}g)$ is given by
(6.7) $a_{x}{\mathfrak{R}}(z_{0}^{2})+a_{y}{\mathfrak{I}}(z_{0}^{2})+a_{\rho}|z_{0}|^{2}+\Phi^{\xi}.$
Moreover, $p$ has one hyperbolic block and the remaining blocks are all elliptic.
By Lemma 2.8, the inclusion $i\colon{\mathbb{C}}\hookrightarrow Y$ sending $z_{0}$ to $[1,0,(z_{0},0,\ldots,0)]$ induces a diffeomorphism $\overline{i}$ between ${\mathbb{C}}/\!/{\mathbb{Z}}_{2}={\mathbb{C}}/{\mathbb{Z}}_{2}$ and ${Y/\!/T}$. By definition of the inclusion $j$ and by (6.7),
$T^{2}_{0}(i^{*}g)(z_{0})=a_{x}{\mathfrak{R}}(z_{0}^{2})+a_{y}{\mathfrak{I}}(z_{0}^{2})+a_{\rho}|z_{0}|^{2},$
and so 00 is a non-degenerate critical point of $i^{*}g$ of index 1. Hence, by the equivariant Morse lemma, there exists an open neighborhood $V$ of $(0,0)\in{\mathbb{R}}^{2}$ and a map $\psi\colon V\to{\mathbb{C}}$ taking $(0,0)$ to 00 that is an equivariant diffeomorphism onto an open set such that $(i\circ\psi)^{*}g(x,y)=xy$ for all $(x,y)\in V$. Let $\pi\colon{\mathbb{R}}^{2}/{\mathbb{Z}}_{2}\to{\mathbb{R}}^{2}$ be the map that sends $[x,y]$ to $(xy,x^{2}-y^{2})$; it is a homeomorphism onto ${\mathbb{R}}^{2}$ and a diffeomorphism on the complement on $(0,0)$. Setting $U:=(\overline{i}\circ\overline{\psi})(V/{\mathbb{Z}}_{2})$ and $\Psi\colon U\to{\mathbb{R}}^{2}$ to be $\pi\circ(\overline{i}\circ\overline{\psi})^{-1},$ the result follows.
Case 2: $\mathbf{P(z)=z_{0}z_{1}}$. In this case, ${\mathbb{C}}^{h+1}\simeq{\mathbb{C}}^{2}\times{\mathbb{C}}^{h-1}$ and $H\simeq S^{1}\times(S^{1})^{h-1}$, where $S^{1}$ acts on ${\mathbb{C}}^{2}$ with weights ±$\pm 1$ and trivially on ${\mathbb{C}}^{h-1}$, and $(S^{1})^{h-1}$ acts on ${\mathbb{C}}^{h-1}$ in the standard fashion and trivially on ${\mathbb{C}}^{2}$. Hence, the span of $\{\Phi_{H}^{\xi}\mid\xi\in{\mathfrak{h}}\}$ equals the span of $\{|z_{k}|^{2}\mid k\geq 2\}$ and $\Phi_{1}(z):=\frac{1}{2}(|z_{0}|^{2}-|z_{1}|^{2})$. Moreover, since the Taylor polynomial $T^{2}_{0}(j^{*}g)$ is $H$-invariant, by Remark 2.12, it lies in the span of $\{{\mathfrak{R}}(z_{0}z_{1}),{\mathfrak{I}}(z_{0}z_{1}),|z_{i}|^{2}\colon i=0,\ldots,h\}$. Therefore, there exist constants $a_{x},a_{y},a_{\rho}\in{\mathbb{R}}$ and $\xi\in{\mathfrak{h}}$ such that $T^{2}_{0}(j^{*}g)$ is given by
(6.8) $a_{x}{\mathfrak{R}}(z_{0}z_{1})+a_{y}{\mathfrak{I}}(z_{0}z_{1})+a_{\rho}(|z_{0}|^{2}+|z_{1}|^{2})+\Phi^{\xi}.$
$a_{x}{\mathfrak{R}}(z_{0}z_{1})+a_{y}{\mathfrak{I}}(z_{0}z_{1})+a_{\rho}(|z_{0}|^{2}+|z_{1}|^{2})+\lambda_{1}\Phi_{1}+\sum\limits_{k=2}^{h}\lambda_{k}|z_{k}|^{2},$
where $\lambda_{j}\in{\mathbb{R}}\smallsetminus\{0\}$ for $j=1,\ldots,h$. The eigenvalues of the corresponding symplectic endomorphism are given by
±$\pm 2\lambda_{1}\sqrt{-1}\pm\sqrt{a_{x}^{2}+a_{y}^{2}-4a_{\rho}^{2}}\,,\,\pm 2\lambda_{2}\sqrt{-1},\,\ldots,\,\pm 2\lambda_{h}\sqrt{-1},$
where the signs in the first term are independent. Hence, since $p$ is non-degenerate and does not have purely elliptic type, Remark 6.1 implies that $a_{x}^{2}+a_{y}^{2}-4a_{\rho}^{2}>0$, that $p$ has one focus-focus block, and that the remaining blocks are elliptic.
Let $i\colon{\mathbb{C}}^{2}\hookrightarrow Y$ be the inclusion sending $(z_{0},z_{1})$ to $[1,0,(z_{0},z_{1},0,\ldots,0)]$. We observe that $i^{*}\omega_{Y}=\omega_{0}$, the standard symplectic form on ${\mathbb{C}}^{2}$; see Remark 4.2. Moreover, since $i({\mathbb{C}}^{2})$ is invariant under the $S^{1}$-action generated by $\Phi_{1}$, the functions $i^{*}\Phi_{1}$ and $i^{*}g$ commute with respect to the induced Poisson bracket. Moreover, by definition of the inclusion $j$ and by (6.8),
$T^{2}_{0}(i^{*}g)(z_{0},z_{1})=a_{x}{\mathfrak{R}}(z_{0}z_{1})+a_{y}{\mathfrak{I}}(z_{0}z_{1})+a_{\rho}(|z_{0}|^{2}+|z_{1}|^{2})+\lambda i^{*}\Phi_{1}(z_{0},z_{1})$
for some $\lambda\in{\mathbb{R}}$. Since $a_{x},a_{y},$ and $a_{\rho}$ are not all zero, the quadratic forms $i^{*}\Phi_{1}$ and $T^{2}_{0}(i^{*}g)$ are linearly independent. Hence the map $(i^{*}\Phi_{1},T^{2}_{0}(i^{*}g))\colon{\mathbb{C}}^{2}\to{\mathbb{R}}^{2}$ is regular on a dense subset. Since the Taylor polynomial of $i^{*}g-T^{2}_{0}(i^{*}g)$ of degree two at zero vanishes, the 2-form $d(i^{*}\Phi_{1})\wedge d(i^{*}g-T^{2}_{0}(i^{*}g))$ goes to 0 more quickly than $d(i^{*}\Phi_{1})\wedge d(T^{2}_{0}(i^{*}g))$ as $z=(z_{0},z_{1})$ approaches 00. Hence, the map $(i^{*}\Phi_{1},i^{*}g)$ is regular on a dense subset of an open neighborhood $W$ of 00 and so $(W,\omega_{0},{\mathbb{R}}\times{\mathbb{R}},(i^{*}\Phi_{1},i^{*}g))$ is an integrable system. Moreover, since $a_{x}^{2}+a_{y}^{2}-4a_{\rho}^{2}>0$, the point $0\in{\mathbb{C}}^{2}$ is a non-degenerate singular point with one focus-focus block. So by Lemma 5.4, there exist an open neighborhood $V$ of $[0]\in{\mathbb{C}}^{2}/\!/S^{1}$ and a map $\psi\colon V\to{\mathbb{R}}^{2}$ taking $[0]$ to $(0,0)$ satisfying (1) – (3). Moreover, by Lemma 2.8, the inclusion $i$ induces a diffeomorphism $\overline{i}\colon{\mathbb{C}}^{2}/\!/S^{1}\to{M/\!/T}$. Setting $U:=\overline{i}(V)$ and $\Psi\colon U\to{\mathbb{R}}^{2}$ to be $\psi\circ\overline{i}^{-1}$, the result follows. ∎
Proposition 6.9.
Let $\left(M,\omega,{\mathfrak{t}}\times{\mathbb{R}},f:=\left(\Phi,g\right)\right)$ be an integrable system such that $\left(M,\omega,\Phi\right)$ is a complexity one $T$-space. Assume that $p\in M$ is tall and non-degenerate. If the defining polynomial of the local model for $p$ has degree $N>2$, then $p$ has purely elliptic type.
Proof.
For simplicity, assume that $g(p)=0$. By the Marle-Guillemin-Sternberg local normal form theorem, we may assume that $M$ is a tall local model $Y=T\times_{H}{\mathfrak{h}}^{\circ}\times{\mathbb{C}}^{h+1}$, that $\Phi=\Phi_{Y}$, and that $p=[1,0,0]$. As in the proof of Lemma 6.3, by Lemma 3.2 we may identify $(T_{p}\mathcal{O})^{\omega}/T_{p}\mathcal{O}$ with ${\mathbb{C}}^{h+1}$ so that the homogeneous moment map for the linearized action at $p$ is the restriction of $(\Phi,T^{2}_{p}g)$ to ${\mathbb{C}}^{h+1}$.
Since $N>2$ and since the Taylor polynomial $T^{2}_{0}(j^{*}g)$ is $H$-invariant, by Remark 2.12, it lies in the span of $\{|z_{i}|^{2}\colon i=0,\dots,h\}$, as does $\Phi_{H}^{\xi}$ for any $\xi\in{\mathfrak{h}}$. Hence, a generic element in the span of $\{T^{2}_{0}(j^{*}g),\Phi_{H}^{\xi}\mid\xi\in{\mathfrak{h}}\}$ can be written as
$\sum\limits_{k=0}^{h}\lambda_{k}|z_{k}|^{2},$
where $\lambda_{k}\in{\mathbb{R}}\smallsetminus\{0\}$ for $k=0,\ldots,h$. The eigenvalues of the corresponding symplectic endomorphism are given by
±$\pm 2\lambda_{0}\sqrt{-1},\,\ldots,\,\pm 2\lambda_{h}\sqrt{-1}.$
Hence, since $p$ is non-degenerate, Remark 6.1 implies that $p$ has purely elliptic type. ∎
Finally, we prove the main result of this section.
Proof of Proposition 6.2.
By Lemma 6.3, every short non-degenerate point has purely elliptic type, we may assume that $p$ is tall. Hence, the result is an immediate consequence of Propositions 6.4, 6.6 and 6.9 below. ∎
@@ 7. Proof of the main result
Let $(M,\omega,{\mathfrak{t}}\times{\mathbb{R}},f=(\Phi,g))$ be an integrable system such that $(M,\omega,\Phi)$ is a complexity one $T$-space. Assume that each tall singular point in $(M,\omega,{\mathfrak{t}}\times{\mathbb{R}},f)$ is non-degenerate. Fix $\alpha\in{\mathfrak{t}}^{*}$. Each component of the reduced space $\Phi^{-1}(\alpha)/T$ containing more than one point can be given the structure of a smooth oriented 2-dimensional manifold such that $\overline{g}\colon\Phi^{-1}(\alpha)/T\to{\mathbb{R}}$ is a Morse function. Moreover, given $p\in\Phi^{-1}(\alpha)$:
@@
    $[p]$ is a critical point of $\overline{g}$ of index $1$ exactly if $p$ has a hyperbolic block and connected $T$-stabilizer, and
@@
    $[p]$ is a critical points of index 00 or $2$ exactly if $p$ has purely elliptic type and is a critical point of $g$ modulo $\Phi$.
In particular, if $p$ has a focus-focus block, has a hyperbolic block with disconnected $T$-stabilizer, or is a regular point of $g$ modulo $\Phi$, then $[p]$ is a regular point of $\overline{g}.$
@@
Assume without loss of generality that $\alpha=0$. For simplicity, assume that $\Phi^{-1}(0)/T$ is connected and contains more than one point. Since the orbit of a short point in $\Phi^{-1}(0)$ is an isolated point in $\Phi^{-1}(0)/T$, every point in $\Phi^{-1}(0)$ is tall. If $p\in\Phi^{-1}(0)$ is a critical point of $g$ modulo $\Phi$, then $p$ is also a singular point of $\left(M,\omega,{\mathfrak{t}}\times{\mathbb{R}},f\right)$. So, by assumption, $p$ is non-degenerate. Therefore, the claim follows immediately from Propositions 3.4, 4.5 and 6.2. ∎
@@
Let $N$ be a connected manifold and let $\overline{g}\colon N\to{\mathbb{R}}$ be a proper Morse-Bott function. If the level set $\overline{g}^{-1}(c)$ is connected for each $c\in{\mathbb{R}}$, then any local extremum is a global extremum.
@@
Let $C$ be a critical manifold that is a local extremum. We may assume that it has index 00. By the Morse lemma, there exists a neighborhood $U$ of $p$ such that $\overline{g}^{-1}(0)\cap U=\{p\}$ and $\overline{g}(U)\subseteq[0,\infty)$. Since the level sets are connected, the former implies that $\overline{g}^{-1}(0)=p$. Hence, the latter implies that $\overline{g}^{-1}[0,\infty)$ is a nonempty subset of $N$ that is both open and closed. Since $N$ is connected, this implies that $\overline{g}$ achieves its minimum at $p$. The argument for the case of coindex zero is entirely analogous. ∎
@@
    the level set $\overline{g}^{-1}(c)$ is connected for all $c\in{\mathbb{R}}$ and $N$ is simply connected.
@@
Assume that no critical point of $\overline{g}$ has index one. Since $N$ is 2-dimensional, by Proposition 5.6 the level set $\overline{g}^{-1}(c)$ is connected for all $c\in{\mathbb{R}}$. Hence, since $N$ is closed, by Lemma 7.2 $\overline{g}$ has exactly two critical points. By [Mil, Theorem 4.1], $N$ is homeomorphic to a sphere and, therefore, simply connected.
Conversely, assume that the level set $\overline{g}^{-1}(c)$ is connected for all $c\in{\mathbb{R}}$ and that $N$ is simply connected. Since $N$ is closed, by Lemma 7.2 $\overline{g}$ has exactly one point of index zero and one point of index two. Hence, by [Mil, Theorem 5.2], the Euler characteristic of $N$ is equal to $2-C_{1}$, where $C_{1}$ is the number of critical points of $\overline{g}$ of index one. Since $N$ is closed and simply connected, it follows that $C_{1}=0$. ∎
@@
Fix $\alpha\in{\mathfrak{t}}^{*}$. Assume without loss of generality that $\alpha=0$. Since $\Phi$ is proper, the reduced space ${M/\!/T}$ is compact and connected by Theorem 5.5.
If the reduced space ${M/\!/T}$ is a single point then obviously it is simply connected and the level set $f^{-1}(0,c)$ is connected for all $c\in{\mathbb{R}}$. Otherwise, by Proposition 7.1, ${M/\!/T}$ is a smooth closed, connected oriented surface so that $\overline{g}\colon{M/\!/T}\to{\mathbb{R}}$ is a Morse function; moreover, given $p\in\Phi^{-1}(0)$, $[p]$ is a critical point of $\overline{g}$ of index $1$ exactly if $p$ has a hyperbolic block and connected $T$-stabilizer. Since $T$ is connected, the level set $f^{-1}(0,c)$ is connected if and only if $\overline{g}^{-1}(c)$ is connected. Hence, the result follows from Lemma 7.3. ∎
To prove Theorem 1.10, we need the following result.
@@
Let $N$ be a closed, connected oriented $2$-dimensional manifold and let $\overline{g}\colon N\to{\mathbb{R}}$ be a Morse function. If the level set $\overline{g}^{-1}(c)$ is connected for each $c\in{\mathbb{R}}$, then each level set contains an even number of critical points of index $1$.
@@
By Lemma 7.2, $N$ has a unique critical point of index 00 (at the minimum), and a unique critical point of index $2$ (at the maximum). Assume that the level set $\overline{g}^{-1}(c)$ contains $k\geq 1$ critical points of index $1$. By the Morse Lemma $c$ is in the interior of $\overline{g}(N)$, and so there are no critical points of index 00 or index $2$ in $\overline{g}^{-1}(c)$. Fix $\epsilon>0$ so that $[c-\epsilon,c+\epsilon]$ contains no other critical values. Since there is only one point of index 00 we know that the preimage $\overline{g}^{-1}((-\infty,c-\epsilon])$ is connected. Hence, the first Betti number of $\overline{g}^{-1}((-\infty,c+\epsilon])$ is the first Betti number of $\overline{g}^{-1}((-\infty,c-\epsilon])$ plus $k$. Since every compact oriented $2$-dimensional manifold with boundary with an odd first Betti number has at least two boundary components, both of these Betti number must be even. Hence $k$ is even. ∎

### v3

@@ 6. Non-degenerate singular points
Let $(Y,\omega_{Y},{\mathfrak{t}}\times{\mathbb{R}},f:=(\Phi_{Y},g))$ be an integrable system, where $Y=T\times_{H}{\mathfrak{h}}^{\circ}\times{\mathbb{C}}^{h+1}$ is a tall local model. Assume that $p=[1,0,0]$ is a non-degenerate critical point of $g$ modulo $\Phi$ with $g(p)=0$. If $j\colon{\mathbb{C}}^{h+1}\to Y$ is the inclusion $z\mapsto[1,0,z]$, then
  • the map ${Y/\!/T}\to{\mathbb{R}}$ taking $t,\alpha,z]\in{Y/\!/T}$ to $T^{1}_{0}(j^{*}g)(z)$ is identically zero, and
  • the zero set of the map ${Y/\!/T}\to{\mathbb{R}}$ taking $t,\alpha,z]\in{Y/\!/T}$ to $T^{2}_{0}(j^{*}g)(z)$ is
    • a single point if $p$ has purely elliptic type,
    • homeomorphic to $\{(x,y)\in{\mathbb{R}}^{2}\mid xy=0\}$ if $p$ has a hyperbolic block and connected $T$-stabilizer, and
    • homeomorphic to $\mathbb{R}$} otherwise.
@@ 7. Proof of the main result
Let $(M,\omega,{\mathfrak{t}}\times{\mathbb{R}},f=(\Phi,g))$ be an integrable system such that $(M,\omega,\Phi)$ is a complexity one $T$-space. Assume that each tall singular point in $(M,\omega,{\mathfrak{t}}\times{\mathbb{R}},f)$ is non-degenerate. Fix $\beta\in{\mathfrak{t}}^{*}$. Each component of the reduced space $\Phi^{-1}(\beta)/T$ containing more than one point can be given the structure of a smooth oriented 2-dimensional manifold such that $\overline{g}\colon\Phi^{-1}(\beta)/T\to{\mathbb{R}}$ is a Morse function. Moreover, given $p\in\Phi^{-1}(\beta)$:
@@
    $[p]$ is a critical point of $\overline{g}$ of index $1$ exactly if $p$ has a hyperbolic block and connected $T$-stabilizer,
@@
    $[p]$ is a critical point of $\overline{g}$ of index 0 or $2$ exactly if $p$ is a critical point of $g$ modulo $\Phi$ with purely elliptic type, and
  3. (3)
    $[p]$ is a regular point of $\overline{g}$ exactly if $p$ has a focus-focus block, has a hyperbolic block with disconnected $T$-stabilizer, or is a regular point of $g$ modulo $\Phi$.
@@
We may assume that $\beta=0$. For simplicity, assume that $\Phi^{-1}(0)/T$ is connected and contains more than one point. Then, by definition, every point in $\Phi^{-1}(0)$ is tall. If $p\in\Phi^{-1}(0)$ is a critical point of $g$ modulo $\Phi$, then $p$ is also a singular point of $\left(M,\omega,{\mathfrak{t}}\times{\mathbb{R}},f\right)$. So, by assumption, $p$ is non-degenerate. Therefore, by Proposition 3.4, the claim follows from Propositions 4.5 and 6.3. ∎
@@
Let $N$ be a connected manifold and let $\overline{g}\colon N\to{\mathbb{R}}$ be a proper Morse function. If the fiber $\overline{g}^{-1}(c)$ is connected for each $c\in{\mathbb{R}}$, then $\overline{g}$ has at most one point of index 0 (or coindex 0).
@@
Let $p$ be a critical point of index 0. We may assume that $\overline{g}(p)=0$. By the Morse lemma, there exists a neighborhood $U$ of $p$ such that $\overline{g}^{-1}(0)\cap U=\{p\}$ and $\overline{g}(U)\subseteq[0,\infty)$. Since the fibers are connected, the former implies that $\overline{g}^{-1}(0)=\{p\}$. Hence, the latter implies that $\overline{g}^{-1}[0,\infty)$ is a nonempty subset of $N$ that is both open and closed. Since $N$ is connected, this implies that $\overline{g}$ achieves its minimum at $p$. Hence, $p$ is the unique point of index 0. ∎
@@
    the fiber $\overline{g}^{-1}(c)$ is connected for all $c\in{\mathbb{R}}$ and $N$ is simply connected.
@@
Assume that no critical point of $\overline{g}$ has index one. Since $N$ is 2-dimensional, by Proposition 5.6 the fiber $\overline{g}^{-1}(c)$ is connected for all $c\in{\mathbb{R}}$. Hence, since $N$ is closed, $\overline{g}$ has exactly two critical points by Lemma 7.2. By [Mil, Theorem 4.1], $N$ is homeomorphic to a sphere and, therefore, simply connected.
Conversely, assume that the fiber $\overline{g}^{-1}(c)$ is connected for all $c\in{\mathbb{R}}$ and that $N$ is simply connected. Since $N$ is closed, $\overline{g}$ has exactly one point of index zero and one point of index two by Lemma 7.2. Hence, by [Mil, Theorem 5.2], the Euler characteristic of $N$ is equal to $2-C_{1}$, where $C_{1}$ is the number of critical points of $\overline{g}$ of index one. Since $N$ is closed and simply connected, it follows that $C_{1}=0$. ∎
@@
Let $\left(M,\omega,{\mathfrak{t}}\times{\mathbb{R}},f:=(\Phi,g)\right)$ be an integrable system such that $(M,\omega,\Phi)$ is a complexity one $T$-space with a proper moment map. Assume that each tall singular point in $\left(M,\omega,{\mathfrak{t}}\times{\mathbb{R}},f\right)$ is non-degenerate and that no such point has a hyperbolic block and connected $T$-stabilizer. Fix $\beta\in{\mathfrak{t}}^{*}$ such that $\Phi^{-1}(\beta)/T$ contains more than one point; we may assume that $\beta=0$. Since $\Phi$ is proper, the reduced space M//T{M/\!/T} is compact and connected by Theorem 5.5. Hence, every point in M//T{M/\!/T} is tall, and so by Proposition 7.1, M//T{M/\!/T} is a smooth closed, connected oriented surface so that $\overline{g}\colon{M/\!/T}\to{\mathbb{R}}$ is a Morse function; moreover, $\overline{g}$ has no critical points of index $1$. Therefore, by Lemma 7.3 the fiber $\overline{g}^{-1}(c)$ is connected for all $c\in{\mathbb{R}}$ and M//T{M/\!/T} is simply connected. Finally, since $T$ is connected, this implies that the fiber $f^{-1}(0,c)$ is connected for all $c\in{\mathbb{R}}$. ∎
To prove Theorem 1.10, we also need the following result.
@@
Let $N$ be a closed, connected oriented $2$-dimensional manifold and let $\overline{g}\colon N\to{\mathbb{R}}$ be a Morse function. If the fiber $\overline{g}^{-1}(c)$ is connected for each $c\in{\mathbb{R}}$, then each fiber contains an even number of critical points of index $1$.
@@
By Lemma 7.2, $N$ has a unique critical point of index 0 (at the minimum), and a unique critical point of index $2$ (at the maximum). Assume that the fiber $\overline{g}^{-1}(c)$ contains $k\geq 1$ critical points of index $1$. By the Morse Lemma $c$ is in the interior of $\overline{g}(N)$, and so there are no critical points of index 0 or index $2$ in $\overline{g}^{-1}(c)$. Fix $\epsilon>0$ so that $c-\epsilon,c+\epsilon]$ contains no other critical values. Since there is only one point of index 0 we know that the preimage $\overline{g}^{-1}((-\infty,c-\epsilon])$ is connected. Hence, the first Betti number of $\overline{g}^{-1}((-\infty,c+\epsilon])$ is the first Betti number of $\overline{g}^{-1}((-\infty,c-\epsilon])$ plus $k$. Since every compact oriented $2$-dimensional manifold with boundary with an odd first Betti number has at least two boundary components, both of these Betti number must be even. Hence $k$ is even. ∎